OpenStudy (anonymous):

What is the 20th term of the sequence that begins -3,6,-12,24....?

OpenStudy (anonymous):

@Jhannybean ?

OpenStudy (rajee_sam):

this is a geometric sequence with a common factor of -2 Nth term of a geometric sequence is given by the formula\[a _{n} = a _{1} r ^{(n-1)}\]

OpenStudy (rajee_sam):

n= no. of the term you want to find and r is the common ratio = -2

OpenStudy (rajee_sam):

n = 20 here

OpenStudy (jhannybean):

This is more of a geometric sequence. You can figure this out if you divide the preceding term by the previous term, so 6/-3 =-2. -12/6 =-2 24/-12 = -2 in a Geometric Sequence, we have ratios instead of "differences (d)" and it follows the equation \[\large a_{n}= a_{1}r^{n-1}\]

OpenStudy (jhannybean):

dammit rajee.....-_- I was explaining to him the difference between two two equations he was confused with, lol.

OpenStudy (jhannybean):

so 20th term would be \[\large a_{20}= (-3)(-2)^{20-1}\] r= -2 and a1 = -3 Now you can simplify it and you'll have your answer

OpenStudy (anonymous):

Im a girl lol and I think I did it wrong.

OpenStudy (jhannybean):

Oh sorry :\ how did you do it?

OpenStudy (anonymous):

I did the () first but idk what to do.. lol

OpenStudy (jhannybean):

the what part first?

OpenStudy (jhannybean):

Remember, follow PEMDAS.

OpenStudy (jhannybean):

\[\large a_{50}=(-3)(-2)^{19}\] you can either find what (-2)^19 equals, or punch the whole thing into your calculator and solve it that way :)

OpenStudy (anonymous):

I divided first or multiply?

OpenStudy (jhannybean):

There is no division xD

OpenStudy (jhannybean):

I'd say multiply the (-2)^19 out FIRST and then MULTIPLY it to the (-3)

OpenStudy (jhannybean):

so... (-3)(-524288)

OpenStudy (anonymous):

1572864?

OpenStudy (jhannybean):

You. are. correct :D

OpenStudy (anonymous):

Thanksss :)

OpenStudy (jhannybean):

No problemo